OpenStudy (anonymous):

What is the area of a circle if the diameter is 12 feet?? Answers: a. 113.0 ft² b. 124.8 ft² c. 37.7 ft² d. 452.2 ft²

OpenStudy (cgreenwade2000):

Area of a circle is A= (pi)*r^2

hartnn (hartnn):

\[A =\pi d^2/4\] just plug in d =12

OpenStudy (anonymous):

Thx @hartnn

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

So its 3.14*144 = 452.16/4???

hartnn (hartnn):

you forgot to divide by 4.

OpenStudy (cgreenwade2000):